# Efficient Reasoning for Large Reasoning Language Models via Certainty-Guided Reflection Suppression

### v1

@@ Abstract
Recent Large Reasoning Language Models (LRLMs) employ long chain-of-thought reasoning with complex reflection behaviors, typically signaled by specific trigger words (e.g., “Wait” and “Alternatively”) to enhance performance. However, these reflection behaviors can lead to the overthinking problem where the generation of redundant reasoning steps that unnecessarily increase token usage, raise inference costs, and reduce practical utility. In this paper, we propose Certainty-Guided Reflection Suppression (CGRS), a novel method that mitigates overthinking in LRLMs while maintaining reasoning accuracy. CGRS operates by dynamically suppressing the model’s generation of reflection triggers when it exhibits high confidence in its current response, thereby preventing redundant reflection cycles without compromising output quality. Our approach is model-agnostic, requires no retraining or architectural modifications, and can be integrated seamlessly with existing autoregressive generation pipelines. Extensive experiments across four reasoning benchmarks (i.e., AIME24, AMC23, MATH500, and GPQA-D) demonstrate CGRS’s effectiveness: it reduces token usage by an average of $18.5\%$ to $41.9\%$ while preserving accuracy. It also achieves the optimal balance between length reduction and performance compared to state-of-the-art baselines. These results hold consistently across model architectures (e.g., DeepSeek-R1-Distill series, QwQ-32B, and Qwen3 family) and scales (4B to 32B parameters), highlighting CGRS’s practical value for efficient reasoning.
@@ 1 Introduction
Large Reasoning Language Models (LRLMs), including OpenAI’s o1/o3 (OpenAI 2024, 2025) and DeepSeek-R1 (Guo et al. 2025), have demonstrated remarkable performance on demanding benchmarks, especially in advanced mathematics and program synthesis. A key factor behind their success is their slow-thinking capability, which combines step-by-step deduction with complex reflection behaviors, including backtracking, exploring alternative strategies, and self-verification of results  (Zhu and Li 2025).
@@
Although these reflection behaviors enable models to self-correct and obtain accurate answers, they often suffer from a significant overthinking problem (Cuadron et al. 2025; Sui et al. 2025; Fan et al. 2025), where LRLMs persist in reasoning even after arriving at correct solutions. This phenomenon leads to substantial increases in token consumption, higher inference costs, and a degraded user experience due to unnecessary delays (Sun et al. 2025; Wang et al. 2025b). In extreme cases, excessively long responses may exceed context window limits, truncating critical information and compromising accuracy (Li et al. 2025a). Addressing this overthinking problem and developing efficient reasoning mechanisms, therefore represents a crucial challenge for improving LRLM performance and practicality.
@@
To address this challenge, training-free methods have recently emerged as lightweight solutions that mitigate overthinking without requiring additional training (Zhu and Li 2025). For instance, prompt-guided approaches like TALE (Han et al. 2025) prompts model to respond within predefined token budgets through carefully designed instructions. While straightforward, these methods rely heavily on the model’s innate instruction-following capabilities. Decoding-manipulation techniques, such as Dynasor (Fu et al. 2024) and DEER (Yang et al. 2025b) dynamically adjust the decoding process to eliminate redundancy by terminating generation when the model reaches sufficient confidence. However, their effectiveness remains highly sensitive to the design of early-exit conditions. Consequently, current methods lack adaptive mechanisms to properly balance reflection suppression with reasoning quality, often either excessively suppressing valid corrections or inadequately curbing redundant reflections.
@@
We propose Certainty-Guided Reflection Suppression (CGRS), a novel method that mitigates overthinking in LRLMs while preserving reasoning accuracy. Previous work has demonstrated that reflection behaviors are typically signaled by specific reflection triggers, i.e., keywords like “Wait”, “Alternatively”, “But”, and “Hmm” (Muennighoff et al. 2025; Zhang et al. 2025; Ding et al. 2025). Our CGRS approach proactively suppresses the model’s generation of these reflection triggers when it exhibits high certainty in its current response, thereby preventing redundant reflection cycles, effectively alleviating overthinking, and preserving reasoning quality.
@@
We evaluate CGRS across diverse open-source LRLMs spanning multiple architectures and scales: the DeepSeek-R1-Distill series (Qwen-7B and Llama-8B) (Guo et al. 2025), QwQ-32B (Team 2025), and the Qwen3 family (4B, 8B, 14B, and 32B) (Yang et al. 2025a). Our evaluation covers four reasoning benchmarks, including mathematical tasks (AIME24 (MAA Committees 2025), AMC23 (AI-MO 2024), and MATH500 (Lightman et al. 2023)) and scientific reasoning (GPQA Diamond (Rein et al. 2024)). As shown in Fig. 1, CGRS achieves significant token reduction ranging from $18.5\%$ to $41.9\%$, while maintaining comparable accuracy. Furthermore, CGRS outperforms existing efficient reasoning approaches, including both prompt-guided and decoding-manipulation methods, by delivering greater token efficiency with minimal accuracy impact. These results solidify CGRS as a superior solution for enabling efficient reasoning in LRLMs.
@@ 2 Related Work
LRLMs have significantly advanced long Chain-of-Thought (CoT) reasoning through supervised fine-tuning and reinforcement learning (Pan et al. 2025). However, prolonged CoT often introduces redundant steps that increase computational costs, a phenomenon known as overthinking (Cuadron et al. 2025; Sui et al. 2025; Fan et al. 2025). To address this issue, many works focus on compressing verbose CoTs into more concise reasoning traces (Sui et al. 2025; Qu et al. 2025; Feng et al. 2025; Wang et al. 2025b; Zhu and Li 2025), which can be broadly categorized into two approaches: training-based methods (Arora and Zanette 2025; Aggarwal and Welleck 2025; Wang et al. 2025a; Ma et al. 2025b; Munkhbat et al. 2025; Chen et al. 2025b; Zhu et al. 2025) and training-free methods (Ma et al. 2025a; Yu, Yu, and Wang 2025; Han et al. 2025; Yang et al. 2025c; Team et al. 2025; Wu et al. 2025; Fu et al. 2024; Jiang et al. 2025; Yang et al. 2025b; Li et al. 2025b; Aytes, Baek, and Hwang 2025; Yang et al. 2025c). While effective, training-based methods often require retraining, careful data curation, and may compromise the general capabilities of the pretrained model. In this work, we focus on training-free methods, which eliminate the need for additional training and can be further divided into four categories: prompt-guided, pipeline-based, model-merging, and decoding-manipulation methods.
@@
This type of method (Ma et al. 2025a; Yu, Yu, and Wang 2025; Han et al. 2025) compress lengthy CoT reasoning through prompt engineering. For instance, TALE (Han et al. 2025) simply prompts the model to solve the problem within token budgets. While simple, these methods heavily depend on the model’s inherent instruction-following capability.
@@
This category of methods (Ong et al. 2025; Aytes, Baek, and Hwang 2025; Yang et al. 2025c) distribute queries or reasoning stages across multiple LRLMs, leveraging the efficiency of small models and the advanced reasoning of larger ones. However, these approaches introduce deployment complexity and additional inference overhead from auxiliary modules.
@@
Such methods (Team et al. 2025; Wu et al. 2025) combine slow- and fast-thinking models via weight averaging to enable long-to-short reasoning. While effective for medium-sized models, this approach underperforms at extreme scales (very large or small models) and lacks precise control over reasoning depth (Zhu and Li 2025).
@@
This type of method (Fu et al. 2024; Jiang et al. 2025; Yang et al. 2025b; Li et al. 2025b) dynamically adjusts the decoding process to eliminate redundancy. For example, Dynasor (Fu et al. 2024) and DEER (Yang et al. 2025b) terminate decoding early when the model reaches sufficient confidence, though their performance depends critically on the exit condition design.
@@ 3.1 Preliminaries
Large reasoning language models demonstrate unique generation patterns during inference. Given a problem $\mathbf{q}$} (with prompt included), the output $\mathbf{o}$} is generated token by token as ${\mathbf{o}}_{t}\sim{\bm{\pi}}(\cdot|{\mathbf{q}},{\mathbf{o}}_{<t})$, where $\bm{\pi}$} is the model. LRLMs organize their output using thinking delimiters (i.e., <think> and </think>), splitting the response into two key parts: slow thinking and conclusion. During slow thinking, LRLMs perform detailed, step-by-step reasoning before summarizing their thought process and delivering the final answer in the conclusion (Zhang et al. 2025).
Within slow thinking, models engage in complex reflection behaviors, including backtracking, exploring alternative approaches, and verifying results. These behaviors are often signaled by specific reflection triggers, i.e., keywords like “Wait”, “Alternatively”, “But”, and “Hmm” (Muennighoff et al. 2025; Zhang et al. 2025; Ding et al. 2025).
Notably, these reflection behaviors cause models to continue reasoning even after reaching an initial answer. They may perform extra validation steps or switch reasoning strategies in later stages, either to double-check previous results or explore different paths. However, this can sometimes lead to unproductive overthinking, where the model enters unnecessary reasoning loops, repeatedly validating or re-evaluating without meaningful progress (Chen et al. 2025a).
@@ 3.2 Certainty Estimation
To estimate the necessity of reflective behaviors during inference, we introduce a certainty score that quantifies the LLM’s confidence in its currently generated response.
@@
where $n=|{\mathbf{a}}|$ is the token length of the probed result, ${\mathbf{p}}_{{\mathbf{a}}_{i}}={\bm{\pi}}(\cdot|{\mathbf{q}},{\mathbf{o}}_{<t},{\mathbf{c}},{\mathbf{a}}_{<i})$ represents the probability distribution of ${\mathbf{a}}_{i}$, $\mathcal{H}({\mathbf{p}})=-{\mathbf{p}}^{\top}\log({\mathbf{p}})$ denotes the token entropy, and $|{\mathbf{V}}|$ is the total number of tokens in the LLM’s vocabulary. The term $\log(|{\mathbf{V}}|)$ represents the maximum possible entropy (i.e., when the probability distribution is uniform across all tokens), serving as a normalization factor. This normalization bounds the certainty score $C$ to $[0,1]$, where a higher $C$ indicates higher LLM confidence (i.e., lower entropy in $\mathbf{a}$}), suggesting a reduced need for subsequent reflection behaviors.
@@ 3.3 Reflection Suppression
Following prior work (Yang et al. 2025c; Fu et al. 2024; Zhang et al. 2025), we consider three trigger categories: (i) core hesitation and transition words (e.g., “But” and “Wait”), (ii) alternative proposal markers (e.g., “Alternatively”), and (iii) colloquial contemplation cues (e.g., “Hmm”), all of which signal potential shifts in the LLM’s reasoning mode. Since a single trigger word has multiple natural language variants (e.g., “wait” as a lowercase variant of “Wait”) that map to different tokens in the vocabulary, we construct a comprehensive trigger token set $S_{\text{trigger}}$ through frequency analysis. Specifically, we identify all possible variants of each reflection trigger in the tokenizer’s vocabulary and then analyze their generation frequencies in reasoning traces. This analysis was performed using reasoning traces from the DeepSeek-R1-Distill-Qwen-7B model (Guo et al. 2025) on both the AIME24 (MAA Committees 2025) and AMC23 (AI-MO 2024) benchmarks. To ensure robustness, we conducted four independent inference runs per dataset, aggregating statistics across all executions. Additional details and the complete list of trigger words and their variants is provided in the Appendix.
@@ 4.1 Experimental Setup
We evaluate on three mathematical reasoning benchmarks: AIME24 (MAA Committees 2025) that contains $30$ high-difficulty problems from the 2024 American Invitational Mathematics Examination, AMC23 (AI-MO 2024) that includes $40$ problems from the 2023 American Mathematics Competitions, and MATH500 (Lightman et al. 2023) that has $500$ multi-step problems covering algebra, geometry, and probability, curated by OpenAI, and one scientific reasoning benchmark: GPQA Diamond (abbreviated as GPQA-D) (Rein et al. 2024), a dataset of $198$ multiple-choice questions on biology, chemistry, and physics, at the post-graduate level.
@@
We conduct experiments on the a series of open-source models spanning different architectures and scales, including DeepSeek-R1-Distill series of models (Qwen-7B and Llama-8B) (Guo et al. 2025), QwQ-32B (Team 2025), and the Qwen3 family (4B, 8B, 14B, and 32B) (Yang et al. 2025a).
@@
All experiments are conducted using the open-source vLLM framework (Kwon et al. 2023) to ensure high-throughput and memory-efficient inference. All decoding uses a temperature of $0.6$ and top-p of $0.95$. Each experiment is repeated three times and the average results are reported. The reflection suppression threshold $\delta$ in our CGRS method is set to $0.9$.
@@
The proposed CGRS method is compared with three types of baselines: (i) Vanilla that performs standard decoding without any intervention, (ii) prompt-guided methods, including NoThinking (Ma et al. 2025a) that prompts the model to bypass intermediate reasoning entirely and generate the final answer directly, and TALE (Han et al. 2025) that prompts the model to solve the problem within token budgets (in our experiments, we set this budget based on the actual token length generated by CGRS), and (iii) decoding-manipulation methods, including Dynasor (Fu et al. 2024) that periodically requests intermediate answers at fixed token intervals and exits early if multiple consecutive answers match and DEER (Yang et al. 2025b) dynamically truncates chain-of-thought generation by detecting high-confidence intermediate answers at transition cues such as “Wait”.

### v2

@@ Abstract
Recent Large Reasoning Language Models (LRLMs) employ long chain-of-thought reasoning with complex reflection behaviors, typically signaled by specific trigger words (e.g., “Wait” and “Alternatively”) to enhance performance. However, these reflection behaviors can lead to the overthinking problem where the generation of redundant reasoning steps that unnecessarily increase token usage, raise inference costs, and reduce practical utility. In this paper, we propose Certainty-Guided Reflection Suppression (CGRS), a novel method that mitigates overthinking in LRLMs while maintaining reasoning accuracy. CGRS operates by dynamically suppressing the model’s generation of reflection triggers when it exhibits high confidence in its current response, thereby preventing redundant reflection cycles without compromising output quality. Our approach is model-agnostic, requires no retraining or architectural modifications, and can be integrated seamlessly with existing autoregressive generation pipelines. Extensive experiments across four reasoning benchmarks (i.e., AIME24, AMC23, MATH500, and GPQA-D) demonstrate CGRS’s effectiveness: it reduces token usage by an average of $18.5\%$ to $41.9\%$ while preserving accuracy and also achieves the optimal balance between length reduction and performance compared to state-of-the-art baselines. These results hold consistently across model architectures (e.g., DeepSeek-R1-Distill series, QwQ-32B, and Qwen3 family) and scales (4B to 32B parameters), highlighting CGRS’s practical value for efficient reasoning.
@@ 1 Introduction
Large Reasoning Language Models (LRLMs), including OpenAI’s o1/o3 (o1; o3) and DeepSeek-R1 (guo2025deepseekr1), have demonstrated remarkable performance on demanding benchmarks, especially in advanced mathematics and program synthesis. A key factor behind their success is their slow-thinking capability, which combines step-by-step deduction with complex reflection behaviors, including backtracking, exploring alternative strategies, and self-verification of results  (zhu2025conciseadaptivethinkinglarge).
@@
Although these reflection behaviors enable models to self-correct and obtain accurate answers, they often suffer from a significant overthinking problem (cuadron2025dangeroverthinkingexaminingreasoningaction; sui2025stopoverthinkingsurveyefficient; fan2025missingpremiseexacerbatesoverthinking), where LRLMs persistently continue reasoning even after arriving at correct solutions. This phenomenon leads to substantial increases in token consumption, higher inference costs, and degraded user experience due to unnecessary delays (sun2025invisibletokensvisiblebills; wang2025harnessingreasoningeconomysurvey). In extreme cases, excessively prolonged responses may exceed context window limits, resulting in truncated critical information and compromised accuracy (li2025thinklesstrainingfreeinferenceefficientmethod). Addressing this overthinking problem and developing efficient reasoning mechanisms, therefore represents a crucial challenge for improving LRLM performance and practicality.
@@
To address this challenge, training-free methods have recently emerged as lightweight solutions that mitigate overthinking without requiring additional training (zhu2025conciseadaptivethinkinglarge). For instance, prompt-guided approaches like TALE (han-etal-2025-token) prompts model responses within predefined token budgets through carefully designed instructions. While straightforward, these methods rely heavily on the model’s innate instruction-following capabilities. Decoding-manipulation techniques, such as Dynasor (fu2025efficientlyscalingllmreasoning) and DEER (yang2025dynamicearlyexitreasoning), dynamically adjust the decoding process to eliminate redundancy by terminating generation when the model reaches sufficient confidence. However, their effectiveness remains highly sensitive to the design of early-exit conditions. Consequently, current methods lack adaptive mechanisms to properly balance reflection suppression with reasoning quality, often either excessively suppressing valid corrections or inadequately curbing redundant reflections.
@@
We propose Certainty-Guided Reflection Suppression (CGRS), a novel method that mitigates overthinking in LRLMs while preserving reasoning accuracy. Previous work has demonstrated that reflection behaviors are typically signaled by specific reflection triggers, i.e., keywords like “Wait”, “Alternatively”, “But”, and “Hmm” (muennighoff2025s1; zhang2025alphaonereasoningmodelsthinking; ding2025thinkingtokenshelptrap). Our CGRS approach proactively suppresses the model’s generation of these reflection triggers when it exhibits high certainty in its current response, thereby preventing redundant reflection cycles and effectively alleviating overthinking while preserving reasoning quality.
@@
We evaluate CGRS across diverse open-source LRLMs spanning multiple architectures and scales: the DeepSeek-R1-Distill series (Qwen-7B and Llama-8B) (guo2025deepseekr1), QwQ-32B (qwq32b), and the Qwen3 family (4B, 8B, 14B, and 32B) (qwen3). Our evaluation covers four reasoning benchmarks, including mathematical tasks (AIME24 (aime), AMC23 (amc23), and MATH500 (math500)) and scientific reasoning (GPQA Diamond (gpqa)). As shown in Fig. 1, CGRS achieves significant token reduction ranging from $18.5\%$ to $41.9\%$, while maintaining comparable accuracy. Furthermore, CGRS outperforms existing efficient reasoning approaches, including both prompt-guided and decoding-manipulation methods, by delivering greater token efficiency with minimal accuracy impact. These results solidify CGRS as a superior solution for enabling efficient reasoning in LRLMs.
@@ 2 Related Work
LRLMs have significantly advanced long Chain-of-Thought (CoT) reasoning through supervised fine-tuning and reinforcement learning (Pan2025ASO). However, prolonged CoT often introduces redundant steps that increase computational costs, a phenomenon known as overthinking (cuadron2025dangeroverthinkingexaminingreasoningaction; sui2025stopoverthinkingsurveyefficient; fan2025missingpremiseexacerbatesoverthinking). To address this issue, many works focus on compressing verbose CoTs into more concise reasoning traces (sui2025stopoverthinkingsurveyefficient; qu2025surveyefficientreasoninglarge; feng2025efficientreasoningmodelssurvey; wang2025harnessingreasoningeconomysurvey; zhu2025conciseadaptivethinkinglarge), which can be broadly categorized into two approaches: training-based methods (arora2025traininglanguagemodelsreason; aggarwal2025l1controllinglongreasoning; wang2025learningthinkinformationtheoreticreinforcement; ma2025cotvalvelengthcompressiblechainofthoughttuning; munkhbat2025selftrainingelicitsconcisereasoning; chen2025reasoninglanguagecomprehensivesurvey; zhu2025surveylatentreasoning) and training-free methods (ma2025reasoningmodelseffectivethinking; yu2025premisescalablestrategicprompt; han-etal-2025-token; yang2025speculativethinkingenhancingsmallmodel; kimiteam2025kimik15scalingreinforcement; wu2025unlockingefficientlongtoshortllm; fu2025efficientlyscalingllmreasoning; jiang2025flashthinkearlyexitmethod; yang2025dynamicearlyexitreasoning; li2025steeringllmthinkingbudget; aytes2025sketchofthoughtefficientllmreasoning; yang2025speculativethinkingenhancingsmallmodel). While effective, training-based methods often require retraining, careful data curation, and may compromise the general capabilities of the pretrained model. In this work, we focus on training-free methods, which eliminate the need for additional training and can be further divided into four categories: prompt-guided, pipeline-based, model-merging, and decoding-manipulation methods.
@@
This type of method (ma2025reasoningmodelseffectivethinking; yu2025premisescalablestrategicprompt; han-etal-2025-token) compress lengthy CoT reasoning through prompt engineering. For instance, TALE (han-etal-2025-token) simply prompts the model to solve the problem within token budgets. While simple, these methods heavily depend on the model’s inherent instruction-following capability.
@@
This category of methods (ong2025routellmlearningroutellms; aytes2025sketchofthoughtefficientllmreasoning; yang2025speculativethinkingenhancingsmallmodel) distribute queries or reasoning stages across multiple LRLMs, leveraging the efficiency of small models and the advanced reasoning of larger ones. However, these approaches introduce deployment complexity and additional inference overhead from auxiliary modules.
@@
Such methods (kimiteam2025kimik15scalingreinforcement; wu2025unlockingefficientlongtoshortllm) combine slow- and fast-thinking models via weight averaging to enable long-to-short reasoning. While effective for medium-sized models, this approach underperforms at extreme scales (very large or small models) and lacks precise control over reasoning depth (zhu2025conciseadaptivethinkinglarge).
@@
This type of method (fu2025efficientlyscalingllmreasoning; jiang2025flashthinkearlyexitmethod; yang2025dynamicearlyexitreasoning; li2025steeringllmthinkingbudget) dynamically adjusts the decoding process to eliminate redundancy. For example, Dynasor (fu2025efficientlyscalingllmreasoning) and DEER (yang2025dynamicearlyexitreasoning) terminate decoding early when the model reaches sufficient confidence, though their performance depends critically on the exit condition design.
@@ 3.1 Preliminaries
Large reasoning language models demonstrate unique generation patterns during inference. Given a problem $\mathbf{q}$} (with prompt included), the output $\mathbf{o}$} is generated token by token as ${\mathbf{o}}_{t}\sim{\bm{\pi}}(\cdot|{\mathbf{q}},{\mathbf{o}}_{<t})$, where $\bm{\pi}$} is the model. LRLMs organize their output using thinking delimiters (i.e., <think> and </think>), splitting the response into two key parts: slow thinking and conclusion. During slow thinking, LRLMs perform detailed, step-by-step reasoning before summarizing their thought process and delivering the final answer in the conclusion (zhang2025alphaonereasoningmodelsthinking).
Within slow thinking, models engage in complex reflection behaviors, including backtracking, exploring alternative approaches, and verifying results. These behaviors are often signaled by specific reflection triggers, i.e., keywords like “Wait”, “Alternatively”, “But”, and “Hmm” (muennighoff2025s1; zhang2025alphaonereasoningmodelsthinking; ding2025thinkingtokenshelptrap).
Notably, these reflection behaviors cause models to continue reasoning even after reaching an initial answer. They may perform extra validation steps or switch reasoning strategies in later stages, either to double-check previous results or explore different paths. However, this can sometimes lead to unproductive overthinking, where the model enters unnecessary reasoning loops, repeatedly validating or re-evaluating without meaningful progress (chen2025think23overthinkingo1like).
@@ 3.2 Certainty Estimation
where $n=|{\mathbf{a}}|$ is the token length of the probed results, ${\mathbf{p}}_{{\mathbf{a}}_{i}}={\bm{\pi}}(\cdot|{\mathbf{q}},{\mathbf{o}}_{<t},{\mathbf{c}},{\mathbf{a}}_{<i})$ represents the probability distribution of ${\mathbf{a}}_{i}$, $\mathcal{H}({\mathbf{p}})=-{\mathbf{p}}^{\top}\log({\mathbf{p}})$ denotes the token entropy, and $|{\mathbf{V}}|$ is the total number of tokens in the LLM’s vocabulary. The term $\log(|{\mathbf{V}}|)$ represents the maximum possible entropy (i.e., when the probability distribution is uniform across all tokens), serving as a normalization factor. This normalization bounds the certainty score $C$ to $[0,1]$, where a higher $C$ indicates higher LLM confidence (i.e., lower entropy in $\mathbf{a}$}), suggesting a reduced need for subsequent reflection behaviors.
@@ 3.3 Reflection Suppression
Following prior work (yang2025speculativethinkingenhancingsmallmodel; fu2025efficientlyscalingllmreasoning; zhang2025alphaonereasoningmodelsthinking), we consider three trigger categories: (i) core hesitation and transition words (e.g., “But” and “Wait”), (ii) alternative proposal markers (e.g., “Alternatively”), and (iii) colloquial contemplation cues (e.g., “Hmm”), all of which signal potential shifts in the LLM’s reasoning mode. Since a single trigger word has multiple natural language variants (e.g., “wait” as a lowercase variant of “Wait”) that map to different tokens in the vocabulary, we construct a comprehensive trigger token set $S_{\text{trigger}}$ through frequency analysis. Specifically, we identify all possible variants of each reflection trigger in the tokenizer’s vocabulary and then analyze their generation frequencies in reasoning traces. This analysis was performed using reasoning traces from the DeepSeek-R1-Distill-Qwen-7B model (guo2025deepseekr1) on both the AIME24 (aime) and AMC23 (amc23) benchmarks. To ensure robustness, we conducted four independent inference runs per dataset, aggregating statistics across all executions. Additional details and the complete list of trigger words and their variants are available in Appendix.
@@ 4 Experiments
Qwen3-32B
Vanilla 67.8 11022 - 95.8 6794 - 93.9 4473 - 71.4 7062 - 82.2 -
TALE 67.8 10688 3.0% 93.3 6533 3.8% 93.6 3857 13.8% 67.0 5916 16.2% 80.4 9.2%
NoThinking 41.1 5635 48.9% 75.0 2221 67.3% 87.0 1054 76.4% 56.6 917 87.0% 64.9 69.9%
Dynasor 64.4 9518 13.6% 92.5 5521 18.7% 85.2 3486 22.1% 58.1 3161 55.2% 75.0 27.4%
DEER 68.9 8697 21.1% 88.2 4878 28.2% 93.2 2533 43.4% 69.6 6310 10.6% 80.0 25.8%
CGRS (ours) 65.6 8128 26.3% 94.2 4766 29.8% 93.1 2993 33.1% 64.0 4535 35.8% 79.2 31.2%
@@ 4.1 Experimental Setup
We evaluate on three mathematical reasoning benchmarks: AIME24 (aime) that contains $30$ high-difficulty problems from the 2024 American Invitational Mathematics Examination, AMC23 (amc23) that includes $40$ problems from the 2023 American Mathematics Competitions, and MATH500 (math500) that has $500$ multi-step problems covering algebra, geometry, and probability, curated by OpenAI, and one scientific reasoning benchmark: GPQA Diamond (abbreviated as GPQA-D) (gpqa), a dataset of $198$ multiple-choice questions on biology, chemistry, and physics, at the post-graduate level.
@@
We conduct experiments on the a series of open-source models spanning different architectures and scales, including DeepSeek-R1-Distill series of models (Qwen-7B and Llama-8B) (guo2025deepseekr1), QwQ-32B (qwq32b), and the Qwen3 family (4B, 8B, 14B, and 32B) (qwen3).
@@
All experiments are conducted using the open-source vLLM framework (vllm) to ensure high-throughput and memory-efficient inference. All decoding use temperature $0.6$ and top-p $0.95$. Each experiment is repeated three times and the average results are reported. The reflection suppression threshold $\delta$ in our CGRS method is set to $0.9$.
@@
The proposed CGRS method is compared with three types of baselines: (i) Vanilla that performs standard decoding without any intervention, (ii) prompt-guided methods, including NoThinking (ma2025reasoningmodelseffectivethinking) that prompts the model to bypass intermediate reasoning entirely and generate the final answer directly, and TALE (han-etal-2025-token) that prompts the model to solve the problem within token budgets (in our experiments, we set this budget based on the actual token length generated by CGRS), and (iii) decoding-manipulation methods, including Dynasor (fu2025efficientlyscalingllmreasoning) that periodically requests intermediate answers at fixed token intervals and exits early if multiple consecutive answers match and DEER (yang2025dynamicearlyexitreasoning) dynamically truncates chain-of-thought generation by detecting high-confidence intermediate answers at transition cues such as “Wait”.